OpenStudy (anonymous):

Use a change of variables to find the following indefinite integral. Check your work by differentiation

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ 3x^3 }{ ? \sqrt{5-6x^4}}dx\]

OpenStudy (anonymous):

omit/ignore the question marks

OpenStudy (primeralph):

let A = 3x^2

OpenStudy (anonymous):

hey e .mork u speak chinese or japanese?

OpenStudy (anonymous):

Japanese

OpenStudy (anonymous):

A nice substitution would be: \[\beta = 5-6x^4 \implies d \beta = -24x^3dx \implies \frac{d \beta}{-8} = 3x^3 dx\] This gives: \[\rightarrow \frac{-1}{8} \int\limits \frac{d \beta}{\sqrt{\beta}}=\frac{-1}{8} \int\limits \beta ^ {- \frac{1}{2}}d \beta\] I'll leave the rest to you. Don't forget to plug what beta is in terms of x after you integrate.

OpenStudy (primeralph):

either way.

OpenStudy (anonymous):

cool, i speak chinese... btw all the extra stuff arrountd the problem is confusing me , glad you dont need help anymore...lol

OpenStudy (anonymous):

I do need help....

OpenStudy (anonymous):

oh well could u re-write the question? maybe i can help you.. idk though

OpenStudy (anonymous):

@malevolence19 is it -1/8 + C?

OpenStudy (anonymous):

-1/4X^(1/2) +C?

OpenStudy (anonymous):

Well when you integrate you'll have: \[\frac{-1}{4} \beta ^{\frac{1}{2}} + C = \frac{-1}{4} \sqrt{5-6x^4} + C\]

OpenStudy (anonymous):

When you differentiate, don't forget the chain rule!

OpenStudy (anonymous):

Okay Thanks soo much!!

OpenStudy (anonymous):

good job malevolence19... Lol of course i geet the question at the end...lol

OpenStudy (anonymous):

good job malevolence19... Lol of course i geet the question at the end...lol

OpenStudy (anonymous):

No offense drizzy, but what are you even talking about?

OpenStudy (anonymous):

nevermind..